# Distribution System Reconfiguration to Mitigate Load Altering Attacks via Stackelberg Games

### v5

@@ IV Mitigating LAA via Reconfiguration
In this section, we introduce a novel technique to mitigate LAAs by reconfiguring the DN topology. We exploit a sequential game-theoretic interaction in which, following the LAA launched by the attacker, the defender reconfigures the network to react optimally to the threat.
IV-A Mitigation Design and Intuition
The intuition behind the proposed defense technique of reconfiguring the DN lies in the analytical insights derived in Section III. Based on our analysis, LAAs targeting the leaf buses of the DN lead to the greatest attack impact (i.e. deviation of the voltage from the nominal values). In this context, reconfiguring the DN, changes the position of the leaf buses, thus alleviating the attack impact. It is worth noting that the proposed mitigation leverages the pre-existing capabilities of the DN (e.g., devices enabling network reconfiguration are primarily installed to reduce power losses and/or voltage deviations) and, hence, does not require new infrastructure.
@@ IV-A Mitigation Design and Intuition
Furthermore, in the proposed scheme, the system will be reconfigured only when an attack is detected, thus avoiding unnecessary reconfigurations to mitigate attacks (note that cyberattacks are somewhat rare events). The attack detection module can be based on existing model-driven or data-driven approaches for detecting LAAs. The reader can refer to past works, including [26] and [27] in this area for more details.
IV-B Stackelberg Game for Attack Mitigation
We model the strategic interaction of the attacker and the defender via a non-cooperative Stackelberg game in which the attacker first chooses a bus to launch an LAA, and subsequently, the defender exploits the flexibility of the system to reconfigure it and mitigate the attack. This sequential game models the reactive mitigation framework. Table II summarizes the assumptions and details of the game.
@@ IV-B Stackelberg Game for Attack Mitigation
Knowledge Potential defensive reactions Grid’s topology
Actions Switching loads Changing the topology
Goal Violate voltage const. Min. attack impact
A Stackelberg game with two players consists of a leader and a follower. The leader always commits their action first to maximize their own objective function by anticipating the follower’s reaction. Then, given the action of the leader, the follower picks their best (i.e., optimal) response to maximize their own objective function. Since LAAs are rare incidents in the network, we propose a reactive mitigation method instead of a preventive one. The proposed Stackelberg game can be defined as =$\mathcal{H}=\{(A,D),(\mathcal{S}_{A},\mathcal{S}_{D}),(F_{A},F_{D})\}$, in which $A$ and $D$ are the players (attacker and defender), $\mathcal{S}_{A},\mathcal{S}_{D}$ denote sets of actions of players, and $F_{A},F_{D}$ denote their rewards.
The attacker’s set of actions, $\mathcal{S}_{A}$, is launching an LAA in any of the buses (one or two buses at a time by assumption). The set of defense actions is the set of all possible system reconfigurations discussed in Section II-D such that $\mathcal{S}_{D}=\{1,2,\ldots,N_{B}\}$ denotes the set of indices of all possible reconfiguration matrices $B=[b_{ij}]_{1\leq i,j\leq N}\in\{0,1\}^{N\times N}$ whose entries meet the connectivity constraints and $N_{B}$ represents the number of such matrices. Henceforth, we denote by $B(d)=[b_{ij}(d)]_{1\leq i,j\leq N}$ the adjacency matrix of the network as a result of a specific defense $d\in\mathcal{S}_{D}$. The attacker aims to maximize the voltage deviation, as a result, we define $F_{A}(d,a)=\sum_{n\in\mathcal{N}}\left|v^{2}_{nom}-v^{2}_{n}(d,a)\right|$, as the attacker’s objective function.
@@
where $pen(d)=\sum_{i=1}^{N}\sum_{j=1}^{N}\left|b^{pre}_{ij}-b_{ij}(d)\right|$ is the penalty term for enforcing a system reconfiguration with the minimum switches possible. Note that $b^{pre}_{ij}$ is the component of the grid’s connectivity matrix before reconfiguration. $F_{D}^{perf}$ is relevant when the defender is capable of perfect attack localization. However, due to noisy measurements, the defender might not be able to do this. Instead, we assume that the defender is only able to locate a neighbourhood of the attack, i.e., a connected cluster of buses which contains the bus under attack. We further assume that the defender has a favourite candidate bus under attack, denoted by $n_{att}$, but does not discard the attack possibilities of other buses in the located cluster. To model this, we define a discrete probability vector $\sigma=[\sigma_{1},\sigma_{2},\ldots,\sigma_{N}]$ with entries:
@@
Taking this uncertainty (of detecting precisely the attack location) into account at the defender’s end results in the following reward:
@@
IV-C Bayesian Optimization
BO provides algorithms for optimizing black-box functions, whose mathematical expression is unknown or too complex to analyze [35]. Instead, BO relies on the function’s observed values for given inputs. This makes BO also suitable for optimizing computationally expensive functions to evaluate exhaustively [36]. In our case, obtaining the best response function for all possible attacks requires solving a separate optimization problem for each attack, leading to significant computational complexity. To address this issue, we consider the attacker’s reward to be a black-box function and utilize BO to explore and approximate the optimal attack efficiently. Similarly to [37], we first build a probabilistic model ($F^{p}_{A}$) for $F_{A}(a,r(a))$ using a Gaussian process $(\mathcal{GP})$. This model is based on sample attacks $\mathcal{T}=\{a^{s}_{i}\}$ and their corresponding $\{F_{A}(a^{s}_{i},r(a^{s}_{i}))\}$, which is obtained using:
@@ IV-C Bayesian Optimization
1:  Sample $\{a^{i}\}_{i=1}^{n_{0}}$ from $\mathcal{S}_{A}$, form the set $\mathcal{T}$, and compute $r(a),\ \forall a\in\mathcal{T}$ by solving       $r(a)=\arg\max_{d\in\mathcal{S}_{D}}F_{D}(d,a)$ s.t. connectivity and power flow constraints as an MILP optimization problem via SCIPY and CVXPY;
22:  Compute $F_{A}(r(a),a),\ \forall a\in\mathcal{T}$;
33:  Construct a probabilistic model of the $F_{A}(r(a),a)$ as in (24);
44:  Find the next query point $(a^{\mathrm{next}})$ using (25), attach it to $\mathcal{T}$, and compute $F_{A}(r(a^{\mathrm{next}}),a^{\mathrm{next}})$;
55:  Repeat until the stop condition is met;
66:  Choose $a^{*}=\arg\max_{a\in\mathcal{T}}F_{A}(r(a),a)$;
77:  Compute $r(a^{*})=\arg\max_{d\in\mathcal{S}_{D}}F_{D}(d,a^{*})$ as in step 1;
@@
The constraints of the reconfiguration optimization (as in steps one and seven of the algorithm 1) are linear in the square of the voltages, hence, justifying our choice of the distance between the squares of voltages in the reward function, leading to a linear program instead of a quadratic one (obtained by a variable change $u_{i}=v_{i}^{2}$). The resulting MILP optimizations are carried out in Python, the optimization modeling language is CVXPY, and the solver is SCIPY. We find the reward of the attacker for $(r(a),a),\ \forall a\in\mathcal{T}$. Finally, the action corresponding to the maximum of $F_{A}(r(a),a)$ is selected as the attacking strategy in our Stackelberg formulation. Additionally, $r(a^{*})$ corresponding to the Stackelberg attack action $a^{*}$ is the optimum defensive strategy at the Stackelberg equilibrium: $d^{*}=r(a^{*})$.
IV-D Resource-Constrained Attacker
@@
To accommodate this feature, we define a new game =$\mathcal{H}^{\prime}=\{(A,D),(\mathcal{S}_{A},\mathcal{S}_{D}),(F^{\prime}_{A},F_{D})\}$ in which the attacker launches the “critical attack” which we call them the “resource-constrained attacker”. Note that the critical attack in each bus is manipulating the minimum devices ($c(n)$) in that bus to cause a voltage constraint violation. Indeed, $c(n)$ for each attack is unique and varies with the attack location. In this modified game, $F^{\prime}_{A}$ comprises two components: the total nodal voltage deviation and the attack magnitude. The attacker seeks to maximize the former while minimizing the latter. However, these two terms cannot be simply summed due to their disparate physical characteristics. Therefore, we propose the following reward:
@@ V Results and Discussions
Our simulations are conducted using the IEEE 33-bus and 69-bus systems. There are 32 and 48 load buses in the test grids, respectively, which could be attacked. The ZIP load coefficients are set to the residential load-type F introduced in [30].
@@ V-A Critical Attack
Time Att. location (bus) 03:00 (Least load) 09:00 18:00 (Peak load)
@@
Since the results in Table III are obtained by the approximation discussed in Section III, we evaluate how effective these attacks are when considering the full AC power flow model. To evaluate the extent of the errors (between the voltages computed using the analytical results and the full AC power flow model), we compare the obtained voltage profile with the results of BFS. Fig. 2 shows the voltage profile calculated by the two methods during the peak load demand and the corresponding critical attack. We remark that the actual errors of computing the nodal voltages with our approximations never exceed $1\%$, thus proving the validity of the analytical results.
@@ V-B LAA Mitigation
Next, we examine the proposed LAA mitigation method. The base load profiles (without LAAs) of 33-bus and 69-bus test cases are $60\%$ and $30\%$ of their nominal load profiles in MATPOWER. For scenarios i) and ii), we consider LAA attacks of magnitude $p_{i}^{a}=q_{i}^{a}=0.30\ p.u.$ in which $i$ is the index of the attacked bus. This magnitude of the attack is significant enough to cause voltage safety violations in all the load buses (except those adjacent to the root) and, hence, needs to be mitigated. In 33-bus grid, {(8-21), (9-15), (12-22), (18-33), (25-29)} and in 69-bus grid {(11-43), (13-21), (15-46), (27-65), (50-59)} are the normally open lines which can be closed during the system reconfiguration. Four scenarios are considered next: i) accurate attack localization, ii) errors in attack localization, iii) errors in attack localization and resource-constrained attacker, and iv) attacks on two nodes.
@@
ii) Errors in Attack Localization: In this scenario, following the worst case accuracy of the detection algorithm in [27], we consider $\rho=0.7$ ($70\%$ of certainty about the location of the attack), and then the remaining $30\%$ is split equally between other buses in the neighbourhood (see Section IV). The results for this scenario are presented in Table IV. In this scenario, the defense action should keep the voltage within the desired constraints, assuming an LAA in any of the candidate buses. This uncertainty causes a system reconfiguration that necessitates more switching. Fig. 4 represents the voltage profile of the 33-bus grid after the reconfiguration if any of the three suspect buses are attacked. We can notice that the attack will be successfully mitigated and the voltage constraint will not be violated in any case.
@@
iii) Resource-Constrained Attacker: Here, we present the resource-constrained attacker introduced in Section IV-D with $\lambda=0.5$ in their objective function. Similar to the two previous cases, the results of this scenario can also be found in Table IV. We observe that in this scenario, the attacker chooses to attack Bus 18, as the impact of load alteration on this bus will be the greatest (as the number of compromised devices is taken explicitly into account). The 69-bus case maintains the same Stackelberg equilibrium actions as the past two scenarios.
@@
V-C Significance of Game-Theoretic Approach
We also consider a non-strategic attacker that does not anticipate the defender’s actions. In this case, first, the attacker launches the attack, which maximizes the total voltage deviation, regardless of potential defensive reactions. Subsequently, the defender solves the optimal reconfiguration problem to mitigate the attack. The results of this approach for the 33-bus test case are presented in Table VI. We remark that both scenarios i) and ii) result in the same output. Compared to Table V for the strategic attacker, the defender always benefits. Indeed, in scenario i), the total voltage deviation is dropped from $0.84$ p.u. to $0.80$ p.u.; while in ii), the number of switches is reduced from $4$ to $2$. Note that mitigation with less switching is preferred (even with slightly higher voltage deviation). Note that the strategic attacker in the 69-bus grid picks the deepest bus (Bus 27) as the victim, which is similar to the non-strategic attacker’s choice.
TABLE VI: Non-strategic attacker’s preferred actions, obligated switching numbers, and total voltage deviations.
i 18 0.80 2
ii 18 0.80 2
To show the efficiency of the approximations introduced in Section II-D, we compare the computation time for the implemented MILP model with that of the similar MISOCP model. Exhibited in Table VII, for the 33-bus grid, a single MISOCP optimization takes $206.20$ s while one MILP optimization requires $19.47$ s, which means a reduction of over $90\%$ in calculation time, while for the bigger system, this reduction is even more significant ($96\%$). To further reduce the computational time, BO is implemented to minimize the number of optimizations required for the Stackelberg equilibrium determination. Table VIII contrasts the optimization counts with and without BO, demonstrating a significant reduction.
TABLE VII: Optimization times in seconds.
Op. type 33-bus grid 69-bus grid
MISOCP 206.20 s 3168.87 s
MILP 19.47 s 103.14 s
TABLE VIII: Number of optimizations required to compute the Stackelberg equillibrium
@@ V-C Significance of Game-Theoretic Approach
VI Conclusions

### v6

@@ IV Mitigating LAA via Reconfiguration
In this section, we introduce a novel technique to mitigate LAAs by reconfiguring the DN topology and adjusting the outputs of DERs. We exploit a sequential game-theoretic interaction in which, following the LAA launched by the attacker, the defender reconfigures the network to react optimally to the threat. But first, we clarify the position of this work in the defense process and cyber-resilience of power grids.
IV-A Defense Steps in Power Grid and Position of This Paper
Cyber resilience in power system involves three stages [27] – (i) pre-attack stage that involves identifying potential attacks and taking preventive measures, (ii) during-attack stage, which includes attack detection and mitigation, and (iii) post-attack stage, which involves locating and isolating the compromised system. In an LAA scenario, a group of devices are controlled by the adversary (e.g., botmaster). As a result, even if the operator is able to locate the attack, it may not be possible to isolate the loads quickly. The proposed method thus falls under the category of attack mitigation (i.e., during-attack stage), which can be viewed as a quick response to the attack to keep the grid’s voltage in the desired range. This provides more time for the cyber-recovery step, which includes exactly locating, isolating, and physically removing the attack (which typically incurs additional delays).
Figure 3: Defense timeline against LAA.
IV-B Mitigation Design and Intuition
The intuition behind the proposed defense technique of reconfiguring the DN lies in the analytical insights derived in Section III. Based on our analysis, LAAs targeting the leaf buses of the DN lead to the greatest attack impact (i.e. deviation of the voltage from the nominal values). In this context, reconfiguring the DN changes the position of the leaf buses, thus alleviating the attack impact. The presence of DERs in the grid challenges the notion that LAAs on leaf nodes always have the most detrimental impact. However, we utilize inverter-based DERs as an additional tool to mitigate the effects of LAAs. In our approach, when inverter-based DERs are present, they adjust their active and reactive power outputs to help maintain an optimal voltage profile across the grid under LAA conditions.
@@ IV-B Mitigation Design and Intuition
It is worth noting that the proposed mitigation leverages the pre-existing capabilities of the DN (e.g., devices enabling network reconfiguration are primarily installed to reduce power losses and/or voltage deviations) and, hence, does not require new infrastructure. Furthermore, in the proposed scheme, the system will be reconfigured only when an attack is detected, thus avoiding unnecessary defensive actions (note that cyberattacks are somewhat rare events). The attack detection module can be based on existing model-driven or data-driven approaches for detecting LAAs. The reader can refer to past works, including [37] and [23] in this area for more details.
IV-C Stackelberg Game for Attack Mitigation
We adopt the common cybersecurity practice of conservatively assuming a strong adversary with full system knowledge. This assumption enables us to evaluate the worst-case impact of cyber-physical attacks on DN operations and DERs. Although this model represents an upper bound in terms of adversarial capability, it is consistent with established practice in the security literature, where strong attackers are assumed to stress-test system vulnerabilities and provide benchmarks for resilience [38, 39, 40].
In practice, the gap between realistic attackers and an omniscient adversary is often smaller than presumed. Due to the high degree of digitization and data availability in modern power systems, an attacker can gain significant visibility into system states. Grid topology, historical demand and generation profiles, market data, and renewable generation forecasts are frequently accessible through public datasets, regulatory filings, or system probing [8]. Moreover, insider threats, compromised DER controllers, and advanced reconnaissance techniques can provide attackers with near-complete knowledge of system conditions.
By adopting a strong attacker model, we ensure that the results capture the maximum achievable disruption and highlight systemic vulnerabilities that may otherwise be overlooked under weaker assumptions. Importantly, if the system is resilient to an omniscient adversary, it will inherently be robust against less capable attackers. Conversely, defensive strategies developed under weaker threat models may fail to provide sufficient protection when adversaries obtain more information or resources than anticipated. Thus, the strong-adversary model is not only a theoretical abstraction but also a practical safeguard, ensuring that mitigation strategies address the full spectrum of plausible threats.
We model the strategic interaction between the attacker and the defender as a non-cooperative Stackelberg game. This framework is the natural mathematical framework modeling an asynchronous, sequential decision process between two opposing and strategic agents [41], where the attacker acts first by choosing an LAA strategy. Subsequently, the defender optimally reacts by reconfiguring the system to mitigate the attack. The non-cooperative framework allows us to consider explicitly the worst-case scenario, which is essential in deriving robust defenses. Table II summarizes the game’s assumptions and details.
In our problem, we assume that complete information about the game is available to both players. Therefore, Bayesian games are not a suitable fit, as they are intended for modeling incomplete information in non-cooperative games [42] and introduce significant complexity. The only uncertainty considered in this work, related to the attack location, is incorporated through a probability distribution directly in the function of the defender, as detailed later in this section. Moreover, since cyberattacks on power grids are relatively infrequent events, we do not consider iterative methods such as reinforcement learning ones, which would require frequent interactions between the attacker and the defender.
@@ IV-C Stackelberg Game for Attack Mitigation
Knowledge Potential defensive reactions, Grid topology Attack strategy, Grid topology
Actions Switching loads Changing the topology & Adjusting DERs output (if any exist)
Goal Max. voltage constraint violation Min. attack impact
A Stackelberg game with two players consists of a leader and a follower. The leader always commits first to maximize their own objective function by anticipating the follower’s reaction. Then, given the action of the leader, the follower picks their best (i.e., optimal) response to maximize their own objective function. Since LAAs are rare incidents in the network, we propose a reactive mitigation method. The proposed Stackelberg game can be defined as $\mathcal{H}=\{(A,D),(\mathcal{S}_{A},\mathcal{S}_{D}),(F_{A},F_{D})\}$, in which $A$ and $D$ are the players (attacker and defender), $\mathcal{S}_{A},\mathcal{S}_{D}$ denote sets of actions of players, and $F_{A},F_{D}$ denote their rewards.
The attacker’s set of actions, $\mathcal{S}_{A}$, is launching LAA in any of the buses (one or two buses at a time by assumption). The set of defense actions is the set of all possible reconfigurations discussed in Section II-D such that $\mathcal{S}_{D}=\{1,2,\ldots,N_{B}\}$ denotes the set of indices of all possible reconfiguration matrices $B=[b_{ij}]_{1\leq i,j\leq N}\in\{0,1\}^{N\times N}$ whose entries meet the connectivity constraints and $N_{B}$ represents the number of such matrices. Henceforth, we denote by $B(d)=[b_{ij}(d)]_{1\leq i,j\leq N}$ the adjacency matrix of the network as a result of a specific defense $d\in\mathcal{S}_{D}$. The attacker aims to maximize the voltage deviation, as a result, we define $F_{A}(d,a)=\sum_{n\in\mathcal{N}}\left|v^{2}_{nom}-v^{2}_{n}(d,a)\right|$, as the attacker’s objective function.
@@
where $pen(d)=\sum_{i=1}^{N}\sum_{j=1}^{N}\left|b^{pre}_{ij}-b_{ij}(d)\right|$ is the penalty term for enforcing a system reconfiguration with the minimum switches possible. $F_{D}^{perf}$ is relevant when the defender is capable of perfect attack localization. However, due to noisy measurements, the defender might not be able to do this. Instead, we assume that the defender is only able to locate a neighborhood of the attack, i.e., a connected cluster of buses which contains the bus under attack. We further assume that the defender has a favourite candidate bus under attack, denoted by $n_{att}$, but does not discard the attack possibilities of other buses in the located cluster. To model this, we define a discrete probability vector $\sigma=[\sigma_{1},\sigma_{2},\ldots,\sigma_{N}]$ with entries:
@@
Taking this uncertainty (of precisely detecting the attack location) into account at the defender’s results in the following reward:
@@
IV-D Bayesian Optimization
BO provides algorithms for optimizing black-box functions, whose mathematical expression is unknown or too complex to analyze [43]. Instead, BO relies on the function’s observed values for given inputs. This also makes BO suitable for optimizing computationally expensive functions that need to be evaluated exhaustively [44]. BO does not have a mathematical guarantee of finding the global optimum, but can significantly reduce the complexity of analyses.
In our case, obtaining the best response function for all possible attacks requires solving a separate optimization problem for each attack, resulting significant computational complexity. To address this issue, we consider the attacker’s reward to be a black-box function and utilize BO to explore and approximate the optimal attack efficiently. Similarly to [45], we first build a probabilistic model ($F^{p}_{A}$) for $F_{A}(a,r(a))$ using a Gaussian process $(\mathcal{GP})$. This model is based on sample attacks $\mathcal{T}=\{a^{s}_{i}\}$ and their corresponding $\{F_{A}(a^{s}_{i},r(a^{s}_{i}))\}$, which is obtained using:
@@ IV-D Bayesian Optimization
[…]
2: Compute $F_{A}(r(a),a),\ \forall a\in\mathcal{T}$;
3: Construct a probabilistic model of the $F_{A}(r(a),a)$ as in (25);
4: Find the next query point $(a^{\mathrm{next}})$ using (26), attach it to $\mathcal{T}$, and compute $F_{A}(r(a^{\mathrm{next}}),a^{\mathrm{next}})$;
5: Repeat until the stop condition is met;
6: Choose $a^{*}=\arg\max_{a\in\mathcal{T}}F_{A}(r(a),a)$;
7: Compute $r(a^{*})=\arg\max_{d\in\mathcal{S}_{D}}F_{D}(d,a^{*})$ as in step 1;
@@
[…]
IV-E Resource-Constrained Attacker
@@
To accommodate this feature, we define a new game $\mathcal{H}^{\prime}=\{(A,D),(\mathcal{S}_{A},\mathcal{S}_{D}),(F^{\prime}_{A},F_{D})\}$ in which the attacker launches the “critical attack” which we call them the “resource-constrained attacker”. Note that the critical attack in each bus is manipulating the minimum devices ($c(n)$) in that bus to cause a voltage constraint violation. Indeed, $c(n)$ for each attack is unique and varies with the attack location. In this modified game, $F^{\prime}_{A}$ comprises two components: the total nodal voltage deviation and the attack magnitude. The attacker seeks to maximize the former while minimizing the latter. However, these two terms cannot be simply summed due to their disparate physical characteristics. Therefore, we propose the following reward:
@@
IV-F Integration of inverter-based DERs
Integration of DERs in DNs is inevitable. As a result, we also include them in our models and deploy them as another resource for alleviating the LAA impact. These resources have a capped generation capacity; however, by adjusting their power factor, the shares of their active and reactive output power can be varied. We utilize this flexibility, in coordination with network reconfiguration, by adding the constraint from (12) (in Subsection II-E) and by replacing (7) with (28).
$\begin{cases}\begin{split}s^{\text{zp}}_{i}(v_{i})=\big[p^{l}_{i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2})-p_{\text{der},i}\big]&\\ +j\big[q^{l}_{i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v^{2}_{i})-q_{\text{der},i}\big],\end{split}&\ \text{if}\ i\in\mathcal{N}^{\text{DER}}\\ s^{\text{zp}}_{i}(v_{i})=p^{l}_{i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2})+jq^{l}_{i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v^{2}_{i}),&\ \text{else}\end{cases}$ (28)
@@ V Results and Discussions
Here, our simulations are conducted using the IEEE 33-bus and 69-bus systems. Fig. 4 presents the base topologies of them, in which dashed lines are normally open (disconnected) lines (i.e., auxiliary links). There are 32 and 48 load buses in the test grids, respectively, which could be attacked. Note that the results are partially extended to the 141 bus grid in subsection V-C. The ZIP load coefficients are set to the residential load-type F introduced in [29].
(a) IEEE 33-bus grid
(b) IEEE 69-bus grid
Figure 4: Base configuration of the test cases
@@ V-A Critical Attack
Device Att. location (bus) 03:00 (Least load) 09:00 18:00 (Peak load)
@@
Since the results in Table III are obtained by the approximation discussed in Section III, we evaluate how effective these attacks are when considering the full AC power flow model. To evaluate the extent of the errors (between the voltages computed using the analytical results and those obtained from the full AC power flow model), we compare the obtained voltage profile with the results of BFS. Fig. 5 shows the voltage profile calculated by the two methods during the peak load demand and the corresponding critical attack. We note that the actual errors in computing the nodal voltages using our approximations never exceed $1\%$, thereby validating the analytical results.
@@ V-B LAA Mitigation
Next, we examine the proposed LAA mitigation method. The base load profiles (without LAAs) of 33-bus and 69-bus test cases are $60\%$ and $30\%$ of their nominal load profiles in MATPOWER. For scenarios i) and ii), we consider LAA attacks of magnitude $p_{i}^{a}=q_{i}^{a}=0.30\ p.u.$ in which $i$ is the index of the attacked bus. The magnitude of the attack is significant enough to cause voltage safety violations in all the load buses (except those adjacent to the root) and, hence, needs to be mitigated. Four scenarios are considered next: i) accurate attack localization, ii) errors in attack localization, iii) errors in attack localization and resource-constrained attacker, and iv) attacks on two nodes.
@@
ii) Errors in Attack Localization: In this scenario, following the worst-case accuracy of the detection algorithm in [23], we consider $\rho=0.7$ ($70\%$ of certainty about the location of the attack), and then the remaining $30\%$ is split equally between other buses in the neighbourhood (see Section IV). The results for this scenario are presented in Table IV. In this scenario, the defense action should keep the voltage within the desired constraints, assuming an LAA in any of the candidate buses. This uncertainty causes a system reconfiguration that necessitates more switching. Fig. 7 represents the voltage profile of the 33-bus grid after the reconfiguration if any of the three suspect buses are attacked. We can notice that the attack will be successfully mitigated and the voltage constraint will not be violated in any case.
@@
iii) Resource-Constrained Attacker: Here, we present the resource-constrained attacker introduced in Section IV-E with $\lambda=0.5$ in their objective function. Similar to the two previous cases, the results of this scenario can also be found in Table IV. We observe that in this scenario, the attacker chooses to attack Bus 18, as the impact of load alteration on this bus will be the greatest (as the number of compromised devices is taken explicitly into account). The 69-bus case maintains the same Stackelberg equilibrium actions as the past two scenarios.
@@
To highlight the contribution of the BO in the reduction of the computational burden of our framework in finding the Stackelberg equilibrium, Table VI contrasts the optimization counts with and without BO, demonstrating a significant reduction.
TABLE VI: Number of optimizations required to compute the Stackelberg equillibrium
@@
V-C 141 Bus Grid
In this subsection, we extend our simulations to the 141-bus grid to show the scalability of our defensive method to bigger networks. The data for the auxiliary links of this grid is obtained from [47]. According to this result, when there is an LAA on bus 79, auxiliary line 2-37 is connected, and line 6-37 is disconnected, which maintains the voltage constraints of this grid with a minimum switching.
V-D DNs with Inverter-based DERs
We extend our model by incorporating inverter-based DERs to obtain more realistic models. It is important to note that adding constraints related to DERs changes the nature of the optimization problem: it is no longer an MILP and must instead be formulated and solved as an MISOCP.
The primary effect of incorporating DERs is an increased resilience of the grid against attacks. In other words, the presence of DERs enhances the grid’s ability to withstand more severe attacks. Fig. 8 illustrates the attack configuration on the 33-bus grid with DERs. The red bus represents the targeted node. Notably, when DERs are present in the grid, leaf nodes are no longer necessarily the most attractive targets for the attacker. Instead, the attack strategy becomes closely correlated with the locations of DERs.
To broaden the analysis, Fig. 9 presents the result for the 69-bus grid with DERs. Fig. 10 shows the Stackelberg equilibrium voltage profiles for both grids, with and without DERs. In both grids, the integration of DERs brings the voltage profile closer to the reference value of 1p.u.
Figure 8: Attack and defense actions in Stackelberg equilibrium for the 33-bus grid with DERs.
Figure 9: Attack and defense actions in Stackelberg equilibrium for the 69-bus grid with DERs.
(a) 33-bus grid
(b) 69-bus grid
Figure 10: Voltages of grids in Stackelberg equilibrium without and with DERs.
V-E Grid with Less Auxiliary Links
While the simulation results shown in this section use standard IEEE test grid topologies (including the data on the auxiliary links), we perform additional simulations to show the effectiveness of the proposed defense on the number of auxiliary links.
For this evaluation, we successively remove three of the four auxiliary links in the 33-bus grid, one at a time. Fig. 11 illustrates the voltage profile of the grid when the LAA is conducted at bus 33 when:
(a) All four auxiliary links are available;
(b) Link 25–29 is unavailable;
(c) Links 25–29 and 8–21 are unavailable;
(d) Links 25–29, 8–21, and 18–33 are unavailable; and,
(e) Links 25–29, 8–21, and 18–33 are unavailable, and the attack magnitude is 20% bigger than (d).
Figure 11: Voltage profile of the 33-bus grid after the LAA on bus 33 and defense when (a) All four auxiliary links are available; (b) Link 25–29 is unavailable; (c) Links 25–29 and 8–21 are unavailable; (d) Links 25–29, 8–21, and 18–33 are unavailable; and (e) Links 25–29, 8–21, and 18–33 are unavailable, and the attack magnitude is 20% bigger than (d).
In all scenarios except (e), the reconfiguration-based defense successfully maintains voltage within the specified limits. However, a decrease in the number of auxiliary links results in a slightly worse voltage profile, as the system has fewer reconfiguration options to limit the attack impact. In particular, in scenario (e), with a larger LAA magnitude and with only one auxiliary link, the attack breaches the voltage limits. This shows that while the proposed defense is effective in most practical cases, further research is required on the topic of optimal number and placement of auxiliary links for cyber defense purposes.
V-F Significance of Game-Theoretic Approach
We also consider a non-strategic attacker that does not anticipate the defender’s actions. In this case, first, the attacker launches the attack, which maximizes the total voltage deviation, regardless of potential defensive reactions. Subsequently, the defender solves the optimal reconfiguration problem to mitigate the attack. The results of this approach for the 33-bus test case are presented in Table VII. We remark that both scenarios i) and ii) result in the same output. Compared to Table V for the strategic attacker, the defender always benefits. Indeed, in scenario i), the total voltage deviation is dropped from $0.84$ p.u. to $0.80$ p.u.; while in ii), the number of switches is reduced from $4$ to $2$. Note that mitigation with less switching is preferred (even with slightly higher voltage deviation). Note that the strategic attacker in the 69-bus grid picks the deepest bus (Bus 27) as the victim, which is similar to the non-strategic attacker’s choice.
TABLE VII: Non-strategic attacker’s preferred actions, obligated switching numbers, and total voltage deviations.
i 18 0.80 2
$i$ 18 0.80 2
To discuss the efficiency of the approximations introduced in Section II, we compare the computation time of the implemented MILP model with that of a similar MISOCP model (for grids without DERs). All computations were performed using the default Intel(R) Xeon(R) CPU @ 2.20GHz processor on Google Colab. As shown in Table VIII, a single MISOCP optimization for the 33-bus grid takes 206 seconds, whereas one MILP optimization requires only 19 seconds. This represents a reduction of over 90% in calculation time, with even more significant reductions observed for larger systems.
TABLE VIII: CPU time required to solve different optimizations.
Op. type 33-bus grid 69-bus grid 141-bus grid
MISOCP 206 s 3169 s $\sim 8$ hrs
MILP 19 s 103 s 1077 s
V-G Real-life Latencies
Switch actuation time in commercial devices could take up to 50ms50~ms [48], while communication latencies can add up to 100ms100~ms [49]. In practice, such reconfigurations can be triggered automatically; however, the grid operator may want to ensure that the control actions do not adversely affect the grid. To this end, we assume system operators can use a command authentication scheme, which simulates the dynamics of the grid to observe the impact of the control action before its execution in the real system [50]. For an operator using modern-day real-time simulators, this verification can add 50ms50~ms [51]. In conclusion, we believe that these actions do not add a significant delay, and the proposed method is still feasible.
VI Conclusions and Future Works